OpenStudy (anonymous):

You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a twelve -inch square as the bottom and you want the slant height to be fourteen inches, how much poster board do you need? Answer two hundred fifty-two square inches five hundred thirty-two square inches four hundred eighty square inches four hundred twenty square inches

OpenStudy (anonymous):

i chose the 4th answer is that correct?

OpenStudy (anonymous):

 Area of one triangle = (1/2)(12)(14) = 84 area of all 4 triangles = 84x4 = 336 area of the base = 12x12 = 144 total area = 336+144 =480

OpenStudy (anonymous):

thanxxxxxxx sooo muchhh!! i have on more question i tried working on it but my answer is not one of the choices

OpenStudy (anonymous):

What is the volume of the prism? http://go.flvs.net/courses/1/flvs_71_3598_d3a973c81406668/ppg/educator_math2/v9/07_TB_13_057.jpg seven hundred twenty-nine cubic yards five hundred sixty-seven cubic yards three hundred seventy-eight cubic yards one thousand one hundred thirty-four cubic

OpenStudy (anonymous):

I do not know why I cannot open that web. Could you plz print screen and attach it here?

OpenStudy (anonymous):

yah one sec

OpenStudy (anonymous):

area of the base = (1/2)(9)(7) = 31.5 volume= base x height = 31.5 x 18 = 567

OpenStudy (anonymous):

Thnxxx! last question' A circle has a diameter of 6 millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth.

OpenStudy (anonymous):

diameter = 6mm circumference = 2 pi diameter = 2(3.14)(6) area =(radius)^2 pi = ((6/2)^2) (3.14) I do not want to find my calculator. I think it is at a dark corner of my room XD Plz calculate the final answer by yourself. Sorry for any inconvenience caused. XD

OpenStudy (anonymous):

yah sure of course thanx for all your help i really appreciate it. if i could i would give you 10 medals!

OpenStudy (anonymous):

the circumference is 18.84

OpenStudy (anonymous):

and the area is 28.26

OpenStudy (anonymous):

Are there calculation mistake?

OpenStudy (anonymous):

no there are not i got the correct answer

OpenStudy (anonymous):

My typing mistake for circumference XD circumference = pi diameter = (3.14)(6)

OpenStudy (anonymous):

it is good that you know I am wrong :D